OpenStudy (bbb911):

Find the Slope if it exists y=3.78-4 now i am not sure what i do with the four

OpenStudy (bbb911):

i mean y=3.78x-4

OpenStudy (rmrjr22):

just m = slope

OpenStudy (bbb911):

so its 3.78?

OpenStudy (rmrjr22):

bingo

OpenStudy (bbb911):

thank you :) and if anybody minds giving me an example of one that isn't just tyring to find the slope cos i got some questions here that have answers that are not slopes

OpenStudy (rmrjr22):

no problem

OpenStudy (luigi0210):

Uhm, examples of equations with no slope?

OpenStudy (bbb911):

something like this 3x + 2y = 6

OpenStudy (rmrjr22):

y = mx + b is a good guide

OpenStudy (rmrjr22):

slope is there

OpenStudy (luigi0210):

That has a slope, just not in the right form

OpenStudy (bbb911):

oh i see you have to solve this one now to get the slope right?

OpenStudy (rmrjr22):

yes

OpenStudy (luigi0210):

Yup!

OpenStudy (rmrjr22):

3x + 2y = 6...... 2y = 6 -3x thats a good start

OpenStudy (bbb911):

i think maybe im trying to get advice hobbs so i am able to solve so i wont' do it wrong hopefully that makes sense

OpenStudy (luigi0210):

Don't be rude hob

OpenStudy (bbb911):

you switched the 2y over right doesn't it have to be negative since you brought it over?

OpenStudy (rmrjr22):

i left it alone

OpenStudy (luigi0210):

actually it's the x that switches

OpenStudy (rmrjr22):

you want y all by itsself so it looks like y = mx + b

OpenStudy (bbb911):

oh oky! so the 2y stays put and the 3 turns to -3x okay imma try and solve this :) one sec

OpenStudy (bbb911):

okay i kinda confused i see how now its 2y=6-3x do i now divde all by 2 or ?

OpenStudy (rmrjr22):

yes

OpenStudy (luigi0210):

Yes

OpenStudy (rmrjr22):

the whole side

OpenStudy (luigi0210):

>.>

OpenStudy (rmrjr22):

y = (6-3x)/2 = ?

OpenStudy (bbb911):

kkk :) 1 sec i feel like im answering my own questions sorry im just really unsure all the time lol

OpenStudy (luigi0210):

It's okay, that's what @rmrjr22 is here for,, helping the people in need :D

OpenStudy (rmrjr22):

i thought thats what @Luigi0210 was here for

OpenStudy (bbb911):

okay is the answer -3/2? haha i thought thts why both rmrjr and luigi are here for :)

OpenStudy (luigi0210):

I'm here being Luigi :P

OpenStudy (bbb911):

lol :)

OpenStudy (bbb911):

so hows my answer?

OpenStudy (luigi0210):

the slope is right.

OpenStudy (bbb911):

thanks guys both of you are the best :)

OpenStudy (luigi0210):

Naw, @rmrjr22 is better

OpenStudy (bbb911):

:) you to :)

OpenStudy (luigi0210):

Pfft ._.